OpenStudy (dragontears1128):

Find the 4th term if the sequend in which a1 = 2 and a n+1 = -4a n + 2.

OpenStudy (dragontears1128):

@AliSoliman

OpenStudy (alisoliman):

it is very easy a 1= 2 a 2 = -4 (a 1) +2 = -6 a 3 = -4 (a 2) +2 = 26 a 4= -4(a 3) +2 = -102